OpenStudy (anonymous):

2x + y = 8 x + y = 4 The lines whose equations are given intersect at A.(4, 0) B.(0, 4) C.all points on the line

OpenStudy (anonymous):

ok i just worked it and got B.

OpenStudy (saifoo.khan):

Lol, its A/

OpenStudy (anonymous):

After you solve it you get x = 4. So y => 2.4 + y = 8 => y = 8-8 = 0 => y = 0

OpenStudy (anonymous):

So it's A (4, 0)